Miranda is measuring the sound intensity level for a speaker by holding a sound level meter. The intensity of the sound wave at her location is 2.6e-07 W/m2.

(a) What is the sound intensity level (SIL) recorded on the meter at Miranda's location? i got the answer of 54.14dB

(b) The volume of the speaker is turned down until the intensity of the sound being produced decreases by a factor of 150. What does Miranda's sound level meter read now? i cant get the right answer in dB
Answer #1

b) for the new intensity

I = 2.6 *10^-7/150

for the sound level

Sound level = 10 * log(2.6 *10^-7/(150 * 10^-12))

Sound level = 32.4 dB

there is one more method using the first answer

new sound level = 54.1 + 10 * log(1/150)

new sound level = 32.4 dB